# Domain-specific Question Answering with Hybrid Search

### v1

@@ Introduction
To determine the optimal chunk size for document segmentation, we experimented with various chunk sizes and overlap values. Table 1 summarizes the NDCG scores obtained for different configurations.
@@
  • Fine-tuned Retriever: This approach uses a dense retriever that has been fine-tuned on our domain-specific dataset, as discussed briefly in the Methods section. The fine-tuning enhances the model’s ability to understand domain-specific terminology and nuances.
@@
Answer Similarity: Measures the degree of correspondence between the generated answer and the golden answer. Groundedness: Evaluates how well the generated answer is supported by the retrieved documents, ensuring factual correctness and context relevance. For each query in the golden dataset, we generated answers using the LLM with contexts retrieved by each strategy. We then prompted GPT-4 to rate the similarity and groundedness of each generated answer compared to the golden answer on a scale from 0 to 1.
@@ Discussion
Our evaluation framework’s inclusion of negative test cases (jailbreak attempts, NSFW queries, and irrelevant queries) demonstrates the system’s resilience to potential misuse. The implemented guardrail mechanism, which checks answer similarity against system prompts, provides an effective defense against jailbreak attempts while maintaining system usability.
@@ Appendix A Model Fine-tuning Details
Appendix B Limitations
While our hybrid search approach demonstrates strong performance for domain-specific question answering, several important limitations should be noted:
Context Awareness Limitations
The current system lacks crucial contextual awareness in several dimensions:
  1. 1.
    User Location Context: The system cannot differentiate between users on different pages or sections of the product (e.g., homepage versus editor), leading to potentially misaligned responses. This becomes particularly evident in answers that reference generic locations like “open homepage” without considering the user’s current position in the product journey.
  2. 2.
    Platform and Device Context: The system does not receive or account for information about:
    • Operating system platform (Windows, iOS, etc.)
    • Device type (mobile phone, desktop computer, tablet)
    • Screen size or resolution
    This limitation can result in instructions that may not be applicable to the user’s specific platform or device configuration.
Language Support
The current implementation is restricted to English-language queries and content. This represents a significant limitation for global enterprise deployment, particularly for multinational organizations requiring multilingual support. The system would need substantial modifications to:
  • Handle queries in multiple languages
  • Process multilingual documentation
  • Generate responses in the user’s preferred language
  • Account for cultural and regional differences in terminology
Multimodal Content Handling
A notable limitation exists in the system’s ability to process and leverage visual content:
  1. 1.
    Visual Context Loss: Many retrieved documents contain valuable screenshots, diagrams, or step-by-step visual guides that demonstrate product functionality. However, these visual elements are currently not:
    • Included in the context provided to the LLM
    • Used to enhance response generation
    • Incorporated into the scoring mechanism
  2. 2.
    Response Limitations: The system cannot:
    • Reference specific visual elements in responses
    • Generate or modify visual content
    • Provide visual feedback or confirmation
This limitation is particularly significant in software documentation, where visual guides often provide crucial information that text alone cannot effectively convey.
Future Work
Addressing these limitations suggests several promising directions for future research:
  1. 1.
    Enhanced Context Integration: Developing mechanisms to capture and utilize user context, including:
    • Real-time user location tracking within the product
    • Platform and device detection
    • Context-aware response generation
  2. 2.
    Multilingual Extension: Expanding the system to support multiple languages through:
    • Multilingual dense retriever training
    • Cross-lingual document retrieval
    • Language-specific response generation
  3. 3.
    Multimodal Enhancement: Incorporating visual content through:
    • Multimodal embedding models
    • Visual content understanding and extraction
    • Integration of visual elements in response generation
    • Screenshot and image-aware context processing
Appendix C Evaluation Metrics Details

### v2

@@ 4 Experiments and Evaluation
To determine the optimal chunk size for document segmentation, we experimented with various chunk sizes and overlap values. Table 1 summarizes the NDCG scores obtained for different configurations (see Appendix B for more details about the metric).
@@
  • Fine-tuned Retriever : As shown in Appendix A this approach uses a dense retriever that has been fine-tuned on our domain-specific dataset, as discussed briefly in the Methods section. The fine-tuning enhances the model’s ability to understand domain-specific terminology and nuances.
@@
Answer Similarity: Measures the degree of correspondence between the generated answer and the golden answer (see Appendix B). Groundedness: Evaluates how well the generated answer is supported by the retrieved documents, ensuring factual correctness and context relevance. For each query in the golden dataset, we generated answers using the LLM with contexts retrieved by each strategy. We then prompted GPT-4 to rate the similarity and groundedness of each generated answer compared to the golden answer on a scale from 0 to 1 (see Appendix B).
@@ Results
Human-GPT Score Correlation Analysis
To validate our LLM-based evaluation approach, we performed a focused correlation analysis using a subset of 14 queries that appeared in both golden dataset and a larger human evaluation dataset (containing approximately 200 queries). For these overlapping queries, we compared GPT-4’s answer accuracy scores with binary human feedback (1 for correct, 0 for incorrect). While this represents a limited sample, it provides valuable initial insights into the alignment between automated and human evaluation methods. Analysis of these scores revealed:
  • Strong positive correlation between GPT-4 and human evaluations (Pearson correlation coefficient = $0.89$)
  • GPT-4’s scores showed good discrimination, with higher scores ($\mu=0.82$) for queries rated correct by humans
  • Only one case of significant disagreement occurred (for the Generative Fill query)
While this analysis is based on a small overlapping set, the strong correlation suggests promise in using LLM-based evaluation as a complement to human assessment. Future work should focus on expanding this comparison using a larger set of consistently evaluated queries to more comprehensively validate the automated evaluation approach.
@@ Discussion
Our evaluation framework’s inclusion of negative test cases (jailbreak attempts, NSFW queries, and irrelevant queries) demonstrates the system’s resilience to potential misuse. Detailed analysis of system performance against potentially problematic queries revealed:
Query Type Performance Details
Jailbreak Attempts 91.7% 11/12 null responses
NSFW Content 100% 6/6 null responses
Irrelevant Queries 100% 12/12 null responses
Table 6: System Performance on Negative Queries
The system demonstrated particularly strong performance in handling NSFW content and irrelevant queries, with perfect accuracy in both categories. For jailbreak attempts, only one query out of twelve received any response, demonstrating robust resilience against prompt injection attempts. The implemented guardrail mechanism, which checks answer similarity against system prompts, provides an effective defense against jailbreak attempts while maintaining system usability.
@@
Future Work
Comprehensive Human Evaluation Framework
Building upon our initial correlation analysis between LLM and human evaluations, future work should prioritize:
  • Development of a standardized evaluation protocol with consistent query sets for both human and automated evaluation, clear criteria for human evaluators and structured feedback collection mechanisms
  • Large-scale human evaluation studies including systematic assessment of system performance across diverse query types, evaluation by both domain experts and end-users, collection of qualitative feedback on answer quality and usefulness and analysis of user satisfaction across different use cases
Enhanced Context Integration
: Developing mechanisms to capture and utilize user context, including:
  • Real-time user location tracking within the product
  • Platform and device detection
  • Context-aware response generation
Multilingual Extension
: Expanding the system to support multiple languages through:
  • Multilingual dense retriever training
  • Cross-lingual document retrieval
  • Language-specific response generation
Multimodal Enhancement
: Incorporating visual content through:
  • Multimodal embedding models
  • Visual content understanding and extraction
  • Integration of visual elements in response generation
  • Screenshot and image-aware context processing
@@ Appendix A Model Fine-tuning Details
Appendix B Evaluation Metrics Details